# Mod $\ell$ non-vanishing of self-dual Hecke $L$-values over CM fields and applications

### v1

@@ 4.1.5. Non-Eisenstein functions: CM case
Then restricting to $X^{\varepsilon}_{U}$ for $\varepsilon\in\{\pm\}$, we define spaces of non-Eisenstein forms ${\rm S}_{2}(U,A)^{\varepsilon}$ and ${\rm S}_{2}(A)^{\varepsilon}$. For $f\in{\rm M}_{2}(U,A)$, recall that $f^{\varepsilon}$ is its restriction to $X^{\varepsilon}_{U}$. We typically take $A$ to be the ring of integers of a finite extension of $\mathbb{Q}$$_{\ell}$ and let $\varpi$ be a uniformizer of $A$.
@@
A form $f\in{\rm M}_{2}(U,A)$ has CM by $K$ if
$T_{v}f=a_{v}f$
for all but finitely many finite places $v$ and $a_{v}$ the Hecke eigenvalue of the theta series associated to a Hecke character $\lambda$ over $K$.
@@
    $f^{\varepsilon}\pmod{\varpi}$ is non-zero.
Then $f^{\varepsilon}\pmod{\varpi}$ is non-Eisenstein.
@@
for $\nu$ a primitive unramified quadratic character of $F_{v}^{\times}$ (cf. [33, Thm. 4.7]). In particular, the $T_{v}$-eigenvalue of a spherical vector is 0.
Now assume that $f^{\varepsilon}\pmod{\varpi}$ is Eisenstein.
By the hypothesis, we have $f^{\varepsilon}(z)\equiv\hskip-10.0pt/\ 0\pmod{\varpi}$ for some $z\in X_{U}^{\varepsilon}$. Then note that
$0\equiv T_{v}f\left(z\begin{pmatrix}1&\\ &\varpi_{v}^{-1}\end{pmatrix}\right)\equiv(q_{v}+1)f^{\varepsilon}(z)\equiv\hskip-10.0pt/\ 0\pmod{\varpi}$,
where the second congruence just follows from the definition $T_{q}=\sum_{i=1}^{q_{v}+1}u_{i}$ for $u_{i}\in{\mathrm{GL}}_{2}(F_{v})$ with ${\rm{N}}(u_{i})=\varpi_{v}$, and the hypothesis that $f^{\varepsilon}\pmod{\varpi}$ is Eisenstein. A contradiction. ∎
@@
Let $U\subset\widehat{B}^{\times}$ be an open subgroup of the form $\prod_{v}U_{v}$. Let $v$ be a place of $F$ unramified in $K/F$ such that $B_{v}$ splits and $U_{v}=U_{0}({\mathfrak{p}}_{v}^{k})_{v}$ for some $k$. For a commutative $\mathbb{Z}$}-algebra AA, let $\beta_{1},\cdots,\beta_{s}\in A$ and ${\mathcal{R}}\in{\mathrm{End}}({\rm M}_{2}(U,A))^{\varepsilon}$ be the endomorphism defined by
@@
Let $f^{\varepsilon}\in{\rm S}_{2}(U,A)^{\varepsilon}$ be so that ${\mathcal{R}}f^{\varepsilon}\in{\rm M}_{2}(A)^{\varepsilon}_{\text{Eis}}$. In the following151515 See also the proof of [11, Lem. 5.5]. we show that $f^{\varepsilon}\in{\rm M}_{2}(U,A)_{\text{Eis}}^{\varepsilon}$.
@@ 4.2. Set-up
Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ with $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$, and $\phi_{\lambda}$ the associated ${\mathrm{GL}}_{2}(F)$-theta series of weight $k=2\Sigma+2\kappa$ and level $U_{0}(\mathfrak{N}_{\lambda})$ for $\mathfrak{N}_{\lambda}:=D_{K/F}{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))$.
@@
(cf. Lemma 3.2). Let $G=B^{\times}$ and $\pi=\pi_{\lambda}$ be the cuspidal automorphic representation of $G({\mathbb{A}})$ associated to $\phi_{\lambda}$ with conductor $\mathfrak{N}_{\lambda}$.
Let ${\mathfrak{p}}\nmid D_{K/F}$ be a finite prime of $F$ and $p$ the rational prime below $\mathfrak{p}$}. Let $\mathfrak{c}$} be an ideal coprime to $p\mathfrak{N}_{\lambda}$. Recall that ${\rm{G}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}$ is the Galois group of ring class extension of conductor ${\mathfrak{c}}{\mathfrak{p}}^{\infty}$. Fix a decomposition
${\rm{G}}_{{\mathfrak{c}}{\mathfrak{p}}^{\infty}}=\Delta\times\Gamma_{{\mathfrak{p}}}$
with $\Delta$ torsion and $\Gamma_{{\mathfrak{p}}}$ torsion-free. Fix a character $\chi_{0}$ of ${\rm{G}}_{\mathfrak{c}}$ that is trivial on $\prod_{v|\mathfrak{N}_{\lambda}^{-},v\nmid{\mathfrak{p}}}K_{v}^{\times}$. Let $\Xi_{\mathfrak{p}}$ be the set of finite order characters of $\Gamma_{{\mathfrak{p}}}$. Let $m_{0}={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\lambda_{{\mathfrak{p}}})$ be the exponential conductor of $\lambda_{\mathfrak{p}}$. For any $\nu\in\Xi_{\mathfrak{p}}$ with ${\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu_{\mathfrak{p}})>m_{0}$, the pair $(\pi_{\lambda},\chi_{0}\nu)$ is self-dual with epsilon factor +1+1 (cf. Proposition 3.1 & Lemma 4.11). Define
@@
Note that $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{\pm}$ and $\Xi_{\lambda\chi_{0}^{-1},{\mathfrak{p}}}^{\pm}$ only differ by finitely many characters.
@@
For ${\mathfrak{p}}\nmid\mathfrak{N}_{\lambda}^{-}$, note that the test vector and the period $\Omega_{\lambda}$ do not depend on $\mathfrak{p}$} (cf. Definition 3.4). Whenever ${\mathfrak{p}}|\mathfrak{N}_{\lambda}^{-}$, we will use the notation $\Omega_{\lambda}^{\{{\mathfrak{p}}\}}$ to emphasise the dependence on $\mathfrak{p}$}.
@@ 4.2.2. Epsilon factors
$L(s,\pi_{\lambda,K}\otimes\chi_{0}\nu)=L(s,\lambda\chi_{0}\nu)L(s,\lambda(\chi_{0}\nu)^{-1}).$
We have the following explicit description of global epsilon factors $\varepsilon(\lambda\chi_{0}\nu)$ and hence that of $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{\pm}$.
@@
Suppose that $\nu\in\Xi_{{\mathfrak{p}}}$ has exponential conductor $n>m_{0}$ at $\mathfrak{p}$}. Then
@@
    Suppose that $v\nmid{\mathfrak{p}}$ and $v$ is unramified in $K/F$. For $w|v$, let $c_{w}$ and $n_{w}$ be the exponential conductor of $\mu_{w}$ and $\psi_{K_{w}/F_{w}}$ respectively. Then
@@
    $=$ $\varepsilon(\mu_{v}^{*},\psi_{K_{v}/F_{v}}).$
    Here the second equality follows from definition of the epsilon factor and the third from the fact that $\nu_{v}|_{F_{v}^{\times}}=1$.
@@
    where $\kappa\in K^{\times}$ such that $\overline{\kappa}=-\kappa$ (cf. [38, Prop. 3.7]).
@@ 4.3.1. Definition
Let $f=\Psi(x_{0}\otimes\varphi)$ be the test vector and $\widehat{\varphi}$ the $\ell$-adic avatar of $\varphi$ as in §3.4. Recall that $\widehat{\varphi}$ is $\ell$-optimal with respect to $x=\rho_{k,\ell}(\varsigma_{\ell})^{-1}x_{0}$ and with values in $O_{\pi,\ell}$. The latter is the $\ell$-adic coefficient ring introduced in §3.4, which we enlarge to also contain the image of $\chi_{0}$. Let $S_{0}$ be the set of prime factors of ${\mathfrak{c}}{\mathfrak{p}}$ as in §3.2.
For $a\in\widehat{K}^{\times}$, let $x_{{\mathfrak{c}},n}(a)=x_{{\mathfrak{c}}{\mathfrak{p}}^{n}}(a)$ be a family of CM points as in §2.2.3. Recall that ${\rm G}_{{\mathfrak{c}},n}={\rm G}_{{\mathfrak{c}}{\mathfrak{p}}^{n}}=K^{\times}\widehat{F}^{\times}\backslash\widehat{K}^{\times}/\widehat{O}_{K,{\mathfrak{c}}{\mathfrak{p}}^{n}}$ is the relative Picard group of ${O}_{K,{\mathfrak{c}}{\mathfrak{p}}^{n}}$. We identify ${\rm G}_{{\mathfrak{c}},n}$ with the Galois group of the ring class field of conductor ${\mathfrak{c}}{\mathfrak{p}}^{n}$ over $K$ via the geometrically normalised reciprocity law. Denote by
$\cdot]_{{\mathfrak{c}},n}:\widehat{K}^{\times}\rightarrow{\rm G}_{{\mathfrak{c}},n},\quad a\mapsto[a]_{{\mathfrak{c}},n}$
the natural projection map. Whenever ${\mathfrak{c}}=1$, we will simply write ${\rm{G}}_{n}$ for ${\rm{G}}_{{\mathfrak{c}},n}$ and $x_{n}(\cdot)$ for $x_{{\mathfrak{c}},n}(\cdot)$.
@@
For integers $n\geq{\mathrm{ord}}_{\mathfrak{p}}({\rm N}_{K/F}({\mathrm{Cond}}(\lambda)))$, the $n$-th theta element associated to the self-dual pair $(f,\chi_{0})$ is given by
$\Theta_{\chi_{0},n}(\pi_{\lambda}):=\displaystyle\displaystyle\sum_{a\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}(a)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(a))}\rangle_{x}\cdot[a]_{{\mathfrak{c}},n}\in O_{\pi,\ell}[{\rm G}_{{\mathfrak{c}},n}].$
@@
For any character $\nu\in\Xi_{\mathfrak{p}}$ of conductor ${\mathfrak{p}}^{n}$, we have
@@
$\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))=$ $\sum_{a\in{\rm G}_{{\mathfrak{c}},n}}\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(a))}\rangle_{x}\cdot\nu\chi_{0}(a)$
$=$ $\sum_{a\in{\rm G}_{{\mathfrak{c}},n}}\langle{\rho_{k,\ell}(a_{\ell}\varsigma_{\ell})x,{\varphi}(x_{{\mathfrak{c}},n}(a))}\rangle_{x}\cdot\nu\chi_{0}(a)$
$=$ $\sum_{a\in{\rm G}_{{\mathfrak{c}},n}}\langle{x_{0},{\varphi}(x_{{\mathfrak{c}},n}(a))}\rangle_{x}\cdot\nu\chi_{0}(a)$
@@
We often fix a Hecke character $\chi_{0}$ over $K$ such that the following conditions hold, which are collectively referred to as the condition (D):
 $\mathrm{cond}(\chi)$ is coprime to $\ell{\mathfrak{p}}\mathfrak{N}_{\lambda}$ and only divisible by places inert in $K/F$, (D)
@@ 4.4.1. Inert case
For $\nu\in\Xi_{\mathfrak{p}}$ with $n:={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\nu_{\mathfrak{p}})>m_{0}:={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\lambda_{{\mathfrak{p}}})$, we have
$\varepsilon(\lambda\chi_{0}\nu)=(-1)^{n-m_{0}}\varepsilon(\lambda\chi_{0})$
by Lemma 4.11. Our main result is the following.
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\pi_{\lambda}$ the associated ${\mathrm{GL}}_{2}(F)$-cuspidal automorphic representation of conductor $\mathfrak{N}_{\lambda}$. Let $\mathfrak{p}$} be a degree one prime of $F$ that is inert in $K$ and $\ell$ a prime such that $(\ell,2{\mathfrak{p}}D_{K/F})=1$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. Then for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\pi_{\lambda,K}\otimes\chi_{0}\nu)}{\Omega_{\lambda}^{\cdot}}\right)=0$
@@
Let $a$ be an integer such that ${\bf k}_{\ell}^{\times}[p^{\infty}]=\mu_{p^{a}}$. Choose a finite extension $\mathcal{O}$} of $\mathbb{Z}$$_{\ell}$ in $\overline{{\mathbb{Z}}}_{\ell}$ so that $\mathcal{O}$} contains $O_{\pi_{\lambda},\ell}$, the image of $\chi_{0}$}, $\mu_{p^{a}}$, and let $\varpi$ be a uniformiser of $\mathcal{O}$}. Put ${\bf k}_{\ell}:={\mathcal{O}}/\varpi{{\mathcal{O}}}$.
Recall that we have an $\ell$-optimal test vector $f$. For the order $R$ in the definition of $f$, let
@@
In particular, $\langle{x,\widehat{\varphi}(g)}\rangle_{x}\pmod{\varpi}\in{\rm M}_{2}(U,{\bf k}_{\ell})$ is a non-zero CM form. So under the condition that $\ell\nmid\prod_{v|{\mathfrak{c}}}(q_{v}^{2}-1)$, as a consequence of Lemma 4.7 and [9, Prop. 3.12], it follows that
$H_{\ell}(g):=\sum_{t\in\prod_{v|{\mathfrak{c}}}O_{K_{v}}^{\times}/O_{K_{v},{\mathfrak{c}}_{v}}^{\times}}\chi_{0}(t)\langle{x,\iota_{\varsigma_{\mathfrak{c}}^{({\mathfrak{c}})}}(t)\widehat{\varphi}(g)}\rangle_{x}\pmod{\varpi}\in{\rm M}_{2}(U^{\prime},{\bf k}_{\ell})$
is also a non-zero CM form, where $U^{\prime}\subset\widehat{B}^{\times}$ is an open compact subgroup such that $U^{\prime({\mathfrak{c}})}=U^{({\mathfrak{c}})}$.
We have
@@
$\Theta_{\chi_{0},n}(\pi_{\lambda})$= $\sum_{a\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}(a)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(a))}\rangle_{x}\cdot[a]_{{\mathfrak{c}},n}\pmod{\varpi}$
$\doteq$ $\sum_{a\in{\rm G}_{n}}\chi_{0}(a)H_{\ell}(x_{{\mathfrak{c}},n}(a))\cdot[a]_{n}\pmod{\varpi}$,

### v2

@@ 4.1.5. Non-Eisenstein functions: CM case
Then restricting to $X^{\varepsilon}_{U}$ for $\varepsilon\in\{\pm\}$, we define spaces of non-Eisenstein forms ${\rm S}_{2}(U,A)^{\varepsilon}$ and ${\rm S}_{2}(A)^{\varepsilon}$. For $f\in{\rm M}_{2}(U,A)$, recall that $f^{\varepsilon}$ is its restriction to $X^{\varepsilon}_{U}$.
We typically take $A$ to be the ring of integers of a finite extension of $\mathbb{Q}$$_{\ell}$. Let $\varpi$ be a uniformiser of $A$ and ${\mathbb{F}}=A/(\varpi)$ the residue field.
@@
A form $f\in{\rm M}_{2}(U,{\mathbb{F}})$ has CM by $K$ if
$T_{v}f=f\cdot a_{v}\pmod{\varpi}$
for all but finitely many finite places $v$, where $a_{v}\in A$ is the Hecke eigenvalue under Hecke operator $T_{v}$ of the theta series associated to a Hecke character $\lambda$ over $K$.
Our mod $\ell$ non-vanishing of Rankin–Selberg $L$-values in the CM case crucially relies on the following non-Eisenstein property.
@@
    $f^{\varepsilon}$ is non-zero for some $\varepsilon\in\{\pm\}$.
Then $f^{\varepsilon}$ is non-Eisenstein.
@@
for $\nu$ a primitive unramified quadratic character of $F_{v}^{\times}$ (cf. [31, Thm. 4.7]). In particular, the $T_{v}$-eigenvalue of a spherical vector is 0.
Now assume that $f^{\varepsilon}$ is Eisenstein.
By the hypothesis, we have $f^{\varepsilon}(z)\neq 0$ for some $z\in X_{U}^{\varepsilon}$. Then note that
$0=T_{v}f\left(z\begin{pmatrix}1&\\ &\varpi_{v}^{-1}\end{pmatrix}\right)=(q_{v}+1)f^{\varepsilon}(z)\neq 0\in{\mathbb{F}},$
where the second congruence just follows from the definition $T_{q}=\sum_{i=1}^{q_{v}+1}u_{i}$ for $u_{i}\in{\mathrm{GL}}_{2}(F_{v})$ with ${\rm{N}}(u_{i})=\varpi_{v}$, and the hypothesis that $f^{\varepsilon}$ is Eisenstein. A contradiction. ∎
The following is a refinement of Lemma 4.7.
@@
Let $U\subset\widehat{B}^{\times}$ be an open subgroup of the form $\prod_{v<\infty}U_{v}$. Let $v$ be a place of $F$ unramified in $K/F$ such that $B_{v}$ splits and $U_{v}=U_{0}({\mathfrak{p}}_{v}^{k})_{v}$ for some $k$. For a commutative ring $A$, let $\beta_{1},\cdots,\beta_{s}\in A$ and ${\mathcal{R}}\in{\mathrm{End}}({\rm M}_{2}(A)^{\varepsilon})$ be the endomorphism defined by
@@
Let $f^{\varepsilon}\in{\rm S}_{2}(U,A)^{\varepsilon}$ be so that ${\mathcal{R}}f^{\varepsilon}\in{\rm M}_{2}(A)^{\varepsilon}_{\text{Eis}}$. In the following [^20] See also the proof of [10, Lem. 5.5]. we show that $f^{\varepsilon}\in{\rm M}_{2}(U,A)_{\text{Eis}}^{\varepsilon}$.
@@ 4.2. Set-up
Let $\Sigma$ be a CM type of $K$. Let $\lambda$ be a self-dual Hecke character over $K$ of infinity type $\Sigma+\kappa(1-c)$ with $\kappa\in{\mathbb{Z}}_{\geq 0}[\Sigma]$, and $\phi_{\lambda}$ the associated ${\mathrm{GL}}_{2}(F)$-theta series, which is a Hilbert modular newform of weight $k=2\Sigma+2\kappa$ and level $U_{0}(\mathfrak{N}_{\lambda})$ for $\mathfrak{N}_{\lambda}:=D_{K/F}{\rm{N}}_{K/F}({\mathrm{Cond}}(\lambda))$.
@@
(cf. Lemma 3.2). Let $G=B^{\times}$ and $\pi=\pi_{\lambda}$ be the cuspidal automorphic representation of $G({\mathbb{A}})$ associated to $\phi_{\lambda}$ with conductor $\mathfrak{N}_{\lambda}$, whose existence follows from the Tunnell–Saito theorem (cf. §3.1.2).
Let $\mathfrak{p}$} be a prime ideal of $F$ and $p$ the rational prime below $\mathfrak{p}$}. Let $\mathfrak{c}$} be an ideal coprime to $p\mathfrak{N}_{\lambda}$. Fix a character $\chi_{0}$ of ${\rm{G}}_{\mathfrak{c}}$ of conductor $\mathfrak{c}$} that is trivial on $\prod_{v|\mathfrak{N}_{\lambda}^{-},v\nmid{\mathfrak{p}}}K_{v}^{\times}$. Let $\Xi_{\mathfrak{p}}$ be the set of finite order characters of $\Gamma_{{\mathfrak{p}}}$. Let $m_{\lambda}={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\lambda_{{\mathfrak{p}}})$ be the exponential conductor of $\lambda_{\mathfrak{p}}$. For any $\nu\in\Xi_{\mathfrak{p}}$ with ${\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu_{\mathfrak{p}})>m_{\lambda}$, the pair $(\pi_{\lambda},\chi_{0}\nu)$ is self-dual with epsilon factor +1+1 (cf. Proposition 3.1 with $S_{0}=\{{\mathfrak{p}}\}\sqcup\{{\mathfrak{q}}\ |\ {\mathfrak{q}}|{\mathfrak{c}}\}$). Define
@@
For ${\mathfrak{p}}\nmid\mathfrak{N}_{\lambda}^{-}$, note that the test vector and the period $\Omega_{\lambda}$ does not depend on $\mathfrak{p}$} (cf. Remark 3.10) . Whenever ${\mathfrak{p}}|\mathfrak{N}_{\lambda}^{-}$, we will use the notation $\Omega_{\lambda}^{\{{\mathfrak{p}}\}}$ to emphasise the dependence on $\mathfrak{p}$}.
In the rest of the section, we will assume:
  • ${\mathfrak{p}}\nmid D_{K/F}$,
  • $(\ell,2{\mathfrak{p}}D_{K/F})=1$.
@@ 4.2.2. Epsilon factors
$L(s,\pi_{\lambda,K}\otimes\chi_{0}\nu)=L(s{+1/2},\lambda\chi_{0}\nu)L(s{+1/2},\lambda(\chi_{0}\nu)^{-1}).$
We have the following explicit description of the global epsilon factor $\varepsilon(\lambda\chi_{0}\nu)$ and hence that of $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{\pm}$.
@@
Suppose that $\nu\in\Xi_{{\mathfrak{p}}}$ has exponential conductor $n>m_{\lambda}$ at $\mathfrak{p}$}. Then we have
@@
    Suppose that $v\nmid{\mathfrak{p}}$ and $v$ is unramified in $K/F$. For $w|v$, let $c_{w}$ and $n_{w}$ be the exponential conductor of $\mu_{w}$ and $\psi_{K_{w}/F_{w}}$ respectively. Then we have
@@
    $=$ $\varepsilon(\mu_{v}^{*},\psi_{K_{v}/F_{v}}),$
    where the second equality follows from definition of the epsilon factor and the third from the fact that $\nu_{v}|_{F_{v}^{\times}}=1$.
@@
    where $\kappa\in K_{\mathfrak{p}}^{\times}$ such that $\overline{\kappa}=-\kappa$ (cf. [36, Prop. 3.7]).
@@ 4.3.1. Definition
Let $f=\Psi(x_{0}\otimes\varphi)$ be the test vector and $\widehat{\varphi}$ the $\ell$-adic avatar of $\varphi$ as in §3.4, where we take $S_{0}={\mathfrak{p}}\sqcup\{{\mathfrak{q}}\ |\ {\mathfrak{q}}|{\mathfrak{c}}\}$ in Definition 3.8. Recall that $\widehat{\varphi}$ is $\ell$-optimal with respect to $x=\rho_{k,\ell}(\varsigma_{\ell})^{-1}x_{0}$ and $\langle{x,\widehat{\varphi}(\cdot)}\rangle_{k-2\Sigma}\in O_{\pi,\ell}$. The latter is the $\ell$-adic coefficient ring introduced in §3.4, which we enlarge to also contain the image of $\chi_{0}$.
[…]
and
[…]
In this section, we will use the canonical identification $r$ and will omit the notation $r$. Denote by
$\cdot]_{{\mathfrak{c}},n}:\widehat{K}^{\times}\rightarrow{\rm G}_{{\mathfrak{c}},n},\quad t\mapsto[t]_{{\mathfrak{c}},n}$
the natural projection map. For $t\in\widehat{K}^{\times}$, let $x_{{\mathfrak{c}},n}(t)=x_{{\mathfrak{c}}{\mathfrak{p}}^{n}}(t)$ be a family of CM points as in §2.4.3.
@@
For integers $n\geq m_{\lambda}$, the $n$-th theta element associated to the self-dual pair $(f,\chi_{0})$ is given by
$\Theta_{\chi_{0},n}(\pi_{\lambda}):=\displaystyle\displaystyle\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}(t)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(t))}\rangle_{{k-2\Sigma}}\cdot[t]_{{\mathfrak{c}},n}\in O_{\pi,\ell}[{\rm G}_{{\mathfrak{c}},n}].$
@@
For any character $\nu\in\Xi_{\mathfrak{p}}$ of exponential conductor $n:={\mathrm{ord}}_{\mathfrak{p}}\mathrm{cond}(\nu)\geq{\mathrm{ord}}_{\mathfrak{p}}\mathfrak{N}_{\lambda}$, we have
@@
where $P(\varsigma^{({\mathfrak{c}}{\mathfrak{p}}^{n})},f,\chi_{0}\nu)$ is the toric period defined as in (3.4).
@@
$\nu(\Theta_{\chi_{0},n}(\pi_{\lambda}))=$ $\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(t))}\rangle_{k-2\Sigma}\cdot\nu\chi_{0}(t)$
$=$ $\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\langle{\rho_{k,\ell}(t_{\ell}\varsigma_{\ell})x,{\varphi}(x_{{\mathfrak{c}},n}(t))}\rangle_{k-2\Sigma}\cdot\nu\chi_{0}(t)$
$=$ $\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\langle{x_{0},{\varphi}(x_{{\mathfrak{c}},n}(t))}\rangle_{k-2\Sigma}\cdot\nu\chi_{0}(t)$
@@
Recall that $\mathfrak{N}_{\lambda}$ is the conductor of $\pi_{\lambda}$. We often fix a Hecke character $\chi_{0}$ of ${\rm G}_{{\mathfrak{c}}}$ such that the following conditions hold, which are collectively referred to as the condition (D):
 ${\mathfrak{c}}=\mathrm{cond}(\chi_{0})$ is coprime to $\ell{\mathfrak{p}}\mathfrak{N}_{\lambda}$ and only divisible by places inert in $K/F$, (D)
@@ 4.4.1. Inert case
For $\nu\in\Xi_{\mathfrak{p}}$ with $n:={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\nu_{\mathfrak{p}})>m_{\lambda}={\mathrm{ord}}_{{\mathfrak{p}}}\mathrm{cond}(\lambda_{{\mathfrak{p}}})$, we have
$\varepsilon(\lambda\chi_{0}\nu)=(-1)^{n-m_{\lambda}}\varepsilon(\lambda\chi_{0})$
by Lemma 4.12. It follows that $\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}=\infty$.
Our main result is the following.
@@
Let $K$ be a CM field and $F$ its maximal totally real subfield. Let $\lambda$ be a self-dual Hecke character over $K$ and $\sigma_{\lambda}$ the associated ${\mathrm{GL}}_{2}(F)$-cuspidal automorphic representation of conductor $\mathfrak{N}_{\lambda}$. Let $\mathfrak{p}$} be a degree one prime of $F$ that is inert in $K$ and $\ell$ a prime such that $(\ell,2{\mathfrak{p}}D_{K/F})=1$. Let $\chi_{0}$ be a finite order anticyclotomic Hecke character over $K$ such that the condition (D) holds. Then for all but finitely many $\nu\in\Xi_{\lambda\chi_{0},{\mathfrak{p}}}^{+}$, we have
${v}_{\ell}\left(\frac{\sqrt{|D_{K}|}\cdot L(1/2,\sigma_{\lambda,K}\otimes\chi_{0}\nu)}{\Omega_{\lambda}^{\cdot}}\right)=0$
@@
Let $p$ be the rational prime below $\mathfrak{p}$}. Choose a finite extension $\mathcal{O}$} of $\mathbb{Z}$$_{\ell}$ in $\overline{{\mathbb{Z}}}_{\ell}$ so that $\mathcal{O}$} contains $O_{\pi_{\lambda},\ell}$, the image of $\chi_{0}$} and $\mu_{p}$, where $\mu_{p^{k}}\subset\overline{{\mathbb{Q}}}^{\times}$ is the group of $p^{k}$-th roots of unity.
Let $s\geq 1$ be the integer such that ${\mathcal{O}}^{\times}[p^{\infty}]=\mu_{p^{s}}\subset\overline{{\mathbb{Q}}}$. Let $\varpi$ be a uniformiser of $\mathcal{O}$} and put ${\bf k}_{\ell}={\mathcal{O}}/\varpi{{\mathcal{O}}}\subset\overline{{\mathbb{F}}}_{\ell}$. Since $\mathfrak{p}$$\nmid\ell$, the reduction map ${\mathcal{O}}\rightarrow{\bf k}_{\ell},x\mapsto x\pmod{\varpi}$ induces an isomorphism
${\mathcal{O}}^{\times}[p^{\infty}]\simeq{\bf k}_{\ell}^{\times}[p^{\infty}]$ (4.7)
of abelian groups.
We have an $\ell$-optimal test vector $f=\Psi(x_{0}\otimes\varphi)$ as in §4.3.1. For the order $R$ in the definition of $f$, let
@@
In particular, $\langle{x,\widehat{\varphi}(g)}\rangle_{k-2\Sigma}\pmod{\varpi}\in{\rm M}_{2}(U,{\bf k}_{\ell})$ is a non-zero CM form.
Fix the identification between relative Picard groups and ring class Galois groups as in (4.5) and (4.6). Recall that we have
@@
$\Theta_{\chi_{0},n}(\pi_{\lambda})=$ $\sum_{t\in{\rm G}_{{\mathfrak{c}},n}}\chi_{0}(t)\langle{x,\widehat{\varphi}(x_{{\mathfrak{c}},n}(t))}\rangle_{k-2\Sigma}\cdot[t]_{{\mathfrak{c}},n}.$